OpenStudy (anonymous):

In a change jar, Tom has nickels and dimes. The number of nickels is two times the number of dimes and the total amount of money in the jar is $1.20. How many dimes are in the jar?

OpenStudy (texaschic101):

.05n + .10d = 1.20 n = 2d now sub 2d in for n in the first equation .05(2d) + .10d = 1.20 0.1d + .10d = 1.20 0.2d = 1.20 d = 1.20/0.2 d = 6 n = 2d n = 2(6) n = 12 check.. .05n + .10d = 1.20 .05(12) + .10(6) = 1.20 .60 + .60 = 1.20 1.20 = 1.20 (correct) there is 6 dimes and 12 nickels

OpenStudy (anonymous):

omg thank you so much